The diver sees the sun at an angle of 50 ° above the horizon. Find the angle at which a fisherman in a boat on the surface of the water sees the Sun, again relative to the horizon. The refractive index of water is 1.33.


1
Expert's answer
2021-03-14T19:13:09-0400


θγ = 90 – 50 = 40º

"n_{air}sinθ_i = n_{water}sinθ_γ \\

1 \times sinθ_i = 1.33 sin40º \\

θ_i = sin^{-1}(1.33 \times sin40º) \\

= 58.75º"

φ = 90 – 58.75 = 31.25º

31.25º is the angle at which a fisherman in a boat on the surface of the water sees the Sun, again relative to the horizon.